Grade 12th passMechanics

If mass of hydrogen atom is 1.7×10-24 gm and interatomic distance in a molecule of hydrogen is 4×10-8 cm, then the moment of inertia of the molecule of hydrogen about the axis passing through the center of mass and perpendicular to the line joining the atoms will be

Ayush Nandanwar
6 Years agoGrade 12th pass

1 Answer

Arun
6 Years ago
Explanation-
# Given-
m = 1.7×10-27 kg
r = 4×10^-10 m
# Solution-
Here,
Total mass M = 2m
Radius of rotation R = r/2
Moment of inertia of hydrogen molecule is given by-
M.I. = MR^2
M.I. = (2m)(r/2)^2
M.I. = mr^2/2
M.I. = 1.7×10-27 × (4×10^-10)^2 / 2
M.I. = 1.36×10^-46 kgm^2